OpenStudy (anonymous):

rationalize the denominator: 8/3+√x

jimthompson5910 (jim_thompson5910):

\[\large \frac{8}{3+\sqrt{x}}\] \[\large \frac{8}{3+\sqrt{x}}\times\frac{3-\sqrt{x}}{3-\sqrt{x}}\] \[\large \frac{8(3-\sqrt{x})}{(3+\sqrt{x})(3-\sqrt{x})}\] \[\large \frac{24-8\sqrt{x}}{9-x}\] So \[\large \frac{8}{3+\sqrt{x}}=\frac{24-8\sqrt{x}}{9-x}\]

OpenStudy (anonymous):

\[8(3-\sqrt(x))/(3+\sqrt(x))(3-\sqrt(x))\] \[24-8\sqrt(x)/(9-x)\]

OpenStudy (anonymous):

when you have a denominator a+b, you multiply both numerator and denominator with a-b. if denominator = a-b, multiply a+b. this is to remove the radical sign below the bar(denominator).